A.   A company has its stock currently selling at $67.50. The company is expected to grow at a constant rate of 7 percent. If the appropriate discount rate is 17 percent, what is the current dividend?

                  A. $6.31                     B.   $6.39               C. $6.75                            D. $5.91                       E. $5.86

B.   A stock will not pay a dividend for 6 years. At the end of the seventh year, it will pay a dividend of $7. This dividend will have constant growth of 4% thereafter. If the required rate of return for this stock is 13%, what is its value today?

                  A. $33.06                               B.   $37.36                         C.   $36.44                          D.   $28.74               E. $30.22

C.   A company recently paid a dividend of $4.00 per share. Management expects dividends to grow at the rate of 8% per year for 3 more years. If the required rate of return on the company's stock is 14%, how much would the stock be worth at the end of three years from today?

             A. $68.02                  B. $71.42                  C.   $82.50                          D. $90.70                   E. $109.81
Answer #1

Value of Stock = Expected Dividend/(Required Return – Growth rate)

67.50 = Expected Dividend/(17%-7%)

Expected Dividend = $6.75

Current Dividend = 6.75/(1.07) = $6.31

i.e. A

Value today = 7/(1.13)7 + 7(1.04)/(13%-4%)(1.13)7
= $37.36

i..e B